OpenStudy (anonymous):

A jar contains 0.25 liter of apple juice and 0.30 liter of grape juice. Melissa poured 0.75 liter of pineapple juice into the jar. She then drank 0.20 liter of the mixture.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Write an expression to represent the total amount of juice left in the jar.

OpenStudy (anonymous):

Identify which property is used in each step to solve the expression.

OpenStudy (texaschic101):

(original amount of juice + added juice) - juice that was drank = juice left in the jar (0.25 + 0.30 + 0.75) - 0.20 = 1.30 - 0.20 = 1.10 left in jar

OpenStudy (texaschic101):

Although I am not sure what property that is

OpenStudy (anonymous):

@countryboy97 do you know what the property this is

OpenStudy (anonymous):

@Jumperman

OpenStudy (anonymous):

I need to know what the property is to texeschic's answer

OpenStudy (jumperman):

I'm not sure. This is new to me.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Awesomeness3291

OpenStudy (anonymous):

@kaylaprincess

OpenStudy (anonymous):

Liters

OpenStudy (anonymous):

I mean like associative property